A firm has the following average cost 𝑨π‘ͺ = 𝟐𝟎𝟎 𝑸 + πŸπ‘Έ βˆ’ πŸ‘πŸ”.


a. Find the stationary point and determine if it is a maximum or a minimum.


b. Find the marginal cost function.


1
Expert's answer
2022-04-07T13:08:29-0400

"AC=200\/Q+2Q-36"

A) Taking the FOC

"AC'=-200Q^{-2}+2=0"

"200=2Q^2"

"100=Q^2"

"Q=-10" "Q=10"

Taking the SOC

"AC''=400Q^{-3}"

"AC''(10)=400(10)^{-3}"

"AC''=0.4>0."

The function is a relative minimum.


B) "MC=\frac{dTC}{dQ}"

"TC=AC\u00d7Q=(200\/Q+2Q-36)Q"

"TC=200+2Q^2-36Q"

"MC=4Q-36"